# Deep Incremental Model Based Reinforcement Learning: A One-Step Lookback Approach for Continuous Robotics Control

### v1

@@ Abstract
Model based reinforcement learning (MBRL) attempts to use an available or a learned model to improve the data efficiency of reinforcement learning. This work proposes a one-step lookback approach that jointly learns the latent-space model and the policy to realize the sample-efficient continuous robotic control, wherein the control-theoretical knowledge is utilized to decrease the model learning difficulty. Specifically, the so-called one-step backward data is utilized to facilitate the incremental evolution model, an alternative structured representation of the robotics evolution model in the MBRL field. The incremental evolution model accurately predicts the robotics movement but with low sample complexity. This is because the formulated incremental evolution model degrades the model learning difficulty into a parametric matrix learning problem, which is especially favourable to high-dimensional robotics applications. The imagined data from the learned incremental evolution model is used to supplement training data to enhance the sample efficiency. Comparative numerical simulations on benchmark continuous robotics control problems are conducted to validate the efficiency of our proposed one-step lookback approach.
Keywords model based reinforcement learning  $\cdot$ incremental evolution model  $\cdot$ continuous robotics control
@@ 1 Introduction
The sampling inefficiency hinders model free reinforcement learning (MFRL) algorithms for continuous robotics control, wherein large amounts of expensive physical interactions between robots and environments are required before learning one satisfactory policy [1]. The motivation to improve the sample efficiency brings the advancement of the model based reinforcement learning (MBRL) field [2]. MBRL is an iterative framework wherein an agent acts to collect data, learns the transition function with the collected data, and leverages the learned model for policy learning. The enhanced sample efficiency is realized by introducing an available or learned model into the policy learning process for generating additional training data [3, 4], computing analytic gradient [5, 6], and serving as the basis of planning for short rollouts[7, 8]. However, the model learning part would inevitably introduce additional computational load. Thereby, one efficient approach to learning a latent space model for MBRL is demanded.
This work utilizes control-theoretical knowledge to offer an alternative approach to parameterizing and learning the model for MBRL. The formulated incremental evolution model hugely decreases the model learning difficulty, and favours generality towards high-dimensional robots.
@@ 2 Related Work
MFRL and MBRL: The MFRL algorithms enjoy generality but suffer sample inefficiency. MBRL relieves the dilemma between generality and efficiency mentioned above via learning a model and using the learned model to enhance sample efficiency. Regarding MBRL, the core problems are how to efficiently learn a latent space model and use the learned model for policy optimization.
@@
Model Learning Methods: The fundamental problem of MBRL is how to efficiently learn a robotics evolution model to boost the data efficiency of the policy learning process. Current works often parameterize the robotics evolution model in the forms such as the local linear time-varying system [9], the Gaussian process parameterized model [6], and the neural network parameterized model [10]. There exist two main branches for learning the associated parameters of the parameterized model. One is a model-accuracy-oriented method [10] and the other is a performance-oriented approach [11]. The widely utilized model-accuracy-oriented methods attempt to learn an as accurate model that accurately predicts the robotics movement, in the same logic as the system identification in the control field. However, the performance-oriented approach argues that an accurate learned model does not necessarily lead to a good performance [11]. The above-mentioned objective mismatch problem is solved by learning a model related to high-performance policy, rather than the one accurately describes the robotics movement. This work proposes an alternative approach to parameterize the model using the one-step backward (OSBK) data. The resulting incremental evolution model serves as a different prediction modality in the MBRL field. The formulated incremental evolution model degrades the model learning problem into an easier parametric matrix learning problem, which is especially favourable for high-dimensional robots.
@@
Model Usage and Policy Learning: When the learned model is ready, the subsequent problem is how to efficiently use the learned model to improve the learning efficiency. The learned model is usually used in a Dyna-style [3, 4] where the learned model is used to generate imagined data for the policy learning process. Note that the imagined data is generated without interacting with the environment. Besides, the learned model is used to compute explicit gradient information [5, 6], which offers inferences for policy learning. The works [7, 8] use the learned model to conduct planning for short rollouts. We follow the Dyna-style to get additional training data for the policy learning process.
3 Background
@@
This paper considers a Markov Decision Process (MDP) represented by $\mathcal{M}=\left(\mathcal{S},\mathcal{A},f,r,\gamma\right)$, where $\mathcal{S}\in\mathbb{R}^{n}$, $\mathcal{A}\in\mathbb{R}^{m}$ are the state and action spaces, respectively; $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function; $\gamma\in\left(0,1\right)$ is the discount factor. Denoting $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$ as the robotics state and action. The robotics movement would be described by the transition function $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ that follows
@@
The continuous function $f$ assumes unknown. The robot in the state $s_{t}$ interacts with its surrounding environment by applying the action $a_{t}$ according to the policy $\pi(a|s)$, and evolves into the next state $s_{t+1}$ and obtains the reward $r(s,a)$. In the following, we show how to learn a latent model $\hat{f}$ in Section 4 and how to use the learned model $\hat{f}$ to improve the learning efficiency in Section 5.
@@ 4 Deep Incremental Model
This section first utilizes the so-called OSBK data and the control-theoretical knowledge to transform the nonlinear transfer function (1) into an equivalent parameterized linear form, whose parameters are then learned by a DNN from input-out data.
@@ 4.1 Incremental Model Construction
This work introduces a constant matrix $L$ to facilitate the following model parameterization. We rewrite the robotic evolution model (1) as
$s_{t+1}=La_{t}+h_{t}$ (2)
@@
$s_{t+1}=s_{t}+(L+\frac{h_{t}-h_{t-1}}{\Delta a_{t}})\Delta a_{t}=s_{t}+L_{t}\Delta a_{t}$ (4)
@@
4.2 Model Parameter Learning
This subsection uses the DNN to learn the parametric matrix $L_{t}$ by minimizing the difference between the true and the predicted next-step values.
This work first represents $L_{t}$ in (4) with a fully connected, multi-layer neural network. Then, the Adam algorithm is adopted to train the NN on the dataset $\mathcal{D}$ via minimizing the model prediction error. Finally, we get the learned incremental evolution model
@@ 4.2 Model Parameter Learning
which is the learned representation of (1). The following section will discuss how to use the learned incremental evolution model (5) to improve the learning efficiency.
The impressive performance of the probabilistic dynamics model ensemble technique in existing MBRL works [7] encourages us to also use our learned deterministic incremental evaluation model (5) in an ensemble way, which helps to relieve model biases.
@@
The data is scarce to learn a satisfactory prediction model at the beginning of the model learning process. We mitigate this problem in practice by choosing an initial value for $L(\theta)$ using the prior knowledge of the evolution model (1) if available. The chosen initial value serves as the learning starting point of the $L(\theta)$ learning process. For example, the prior mass matrix of the robot manipulator, although not accurate, is enough to act as the initial value of the $L(\theta)$ learning process.
Figure 1: Schematic of incremental model based reinforcement learning
@@ 5 Policy Learning with Learned Model
This section presents how to use the learned model to improve learning efficiency and how to get the policy for continuous robotics control.
@@ 5.1 Model Usage
This subsection uses the learned incremental evolution model (5) to conduct one-step forward prediction. These predictions and the collected environmental data together serve as sample data for policy evaluation and improvement.
@@ 5.2 Policy Optimization
This work focuses on the continuous robotics control problem. Thereby, we choose the soft actor-critic (SAC) algorithm, a widely accepted algorithm, to solve our problem. In SAC, the critic agent estimates
$Q^{\pi}(s,a)=E_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$ (6)
using the Bellman backup operator,
and the actor finds the policy $\pi$ via minimizing the expected KL-divergence
$J_{\pi}(\phi,\mathcal{D})=\mathbb{E}_{s_{t}\sim\mathcal{D}}\left[D_{KL}(\pi||\exp\left\{Q^{\pi}-V^{\pi}\right\})\right]$ (7)
@@
Note that our learned incremental evolution model (5) is agnostic to the policy learning algorithm. This implies the generality of the MBRL framework in this work. The model learning and usage modules could be used together with different policy optimization algorithms.
Algorithm 1 Incremental Model Based Reinforcement Learning
@@
3:     Train model $f(\theta)$ with $\mathcal{D}_{\text{env}}$ via maximum likelihood
@@
8:              Perform k-step model rollout starting from the sample using policy and add to $\mathcal{D}_{\text{model}}$
@@ 6 Numerical Simulation
The MBPO work inspires this work. This work follows the MBPO work’s way to use the learned model and to optimize a policy. Through the comparison with this milestone work, we check whether our modelling technique is more efficient than other modelling techniques. The comparison results presented in Figure 3 show that our approach achieves comparable performance with the MBPO. However, the modelling technique in our approach holds the potential to apply to high-dimensional robots, while the Gaussian process modelling technique used in MBPO is restricted to the low-dimensional domain.
@@ 7 Discussion
This work jointly learns and improves the model and policy using the interaction with the environment. The control-theoretical knowledge and the OSBK data represent the evolution model of robotics as one linear incremental form that contributes to efficient model learning. The learned incremental evolution model serves as the prediction model in MBRL and improves the sampling efficiency. The formulated incremental evolution model serves as one promising alternative modeling technique in MBRL. The comparative numerical simulations on benchmark continuous robotics control tasks validate the efficiency of our approach.
Limitations The limitation is that how much nonlinearity the incremental evolution model could be addressed remains to be clarified. Besides, the incremental evolution model is hard to represent discontinuous dynamic systems.
Acknowledgements We thank Bruno for the insightful discussions.
@@ 8 Future Works
The utilized control-theoretical knowledge benefits RL with enhanced sample efficiency by offering an explicit control-oriented incremental evolution model, rather than just a black-box input-output mapping. Given the stability analysis and the safety check both require one mathematical form of the controlled plant, the explicit learned model offers us avenues to address safety and stability concerns, which remain to be further explored. This work uses the learned model to generate additional training data. This is one of the ways to utilize the learned model. It is interesting to explore the different roles of the learned model in the learning process. For example, the learned incremental evolution model serves as a basis for planning as in the work .

### v2

# Deep Incremental Model Informed Reinforcement Learning for Continuous Robotic Control

@@ Abstract
Model-based reinforcement learning attempts to use an available or learned model to improve the data efficiency of reinforcement learning. This work proposes a one-step lookback approach that jointly learns the deep incremental model and the policy to realize the sample-efficient continuous robotic control, wherein the control-theoretical knowledge is utilized to decrease the model learning difficulty and facilitate efficient training. Specifically, we use one-step backward data to facilitate the deep incremental model, an alternative structured representation of the robotic evolution model, that accurately predicts the robotic movement but with low sample complexity. This is because the formulated deep incremental model degrades the model learning difficulty into a parametric matrix learning problem, which is especially favourable to high-dimensional robotic applications. The imagined data from the learned deep incremental model is used to supplement training data to enhance the sample efficiency. Comparative numerical simulations on benchmark continuous robotics control problems are conducted to validate the efficiency of our proposed one-step lookback approach.
Keywords Model based reinforcement learning  $\cdot$ deep incremental model  $\cdot$ continuous robotics control
@@ 1 Introduction
The sampling inefficiency hinders model free reinforcement learning (MFRL) algorithms for continuous robotics control, wherein large amounts of expensive physical interactions between robots and environments are required before learning one satisfactory policy [1]. The motivation to improve the sample efficiency brings the advancement of the model based reinforcement learning (MBRL) field [2]. MBRL is an iterative framework wherein an agent acts to collect data, learns the transition function with the collected data, and leverages the learned model for policy learning. The enhanced sample efficiency is realized by introducing an available or learned model into the policy learning process for generating additional training data [3, 4], computing analytic gradient [5, 6], and serving as the basis of planning for short rollouts [7, 8]. However, the model learning part would inevitably introduces additional computational load. Therefore, one efficient approach to learning a latent-space model for MBRL is highly demanded.
This work utilizes control-theoretical knowledge to offer an alternative approach to parameterizing and learning the latent-space model for MBRL. The formulated deep incremental model hugely decreases the model learning difficulty, and favours generality towards high-dimensional robots.
@@ 2 Related Work
MFRL and MBRL: The MFRL algorithms enjoy generality but suffer sample inefficiency. MBRL relieves the dilemma between generality and efficiency mentioned above via learning a model and using the learned model to enhance sample efficiency. Regarding MBRL, the core problems are how to efficiently learn a latent-space model and use the learned model for policy optimization.
@@
Model learning methods: The fundamental problem of MBRL is how to efficiently learn a robotics evolution model to boost the data efficiency of the policy learning process. Current works often parameterize the robotics evolution model in forms such as local linear time-varying system [9], Gaussian process parameterized model [6], and neural network parameterized model [10]. There exist two main branches for learning the associated parameters of the parameterized model. One is the model-accuracy-oriented method [10] and the other is the performance-oriented approach [11]. The widely utilized model-accuracy-oriented methods attempt to learn an accurate model that precisely predicts the robotic movement, in the same logic as the system identification in the control field. However, the performance-oriented approach argues that an accurate learned model does not necessarily lead to a good performance [11]. The above-mentioned objective mismatch problem is solved by learning a model related to high-performance policy, rather than the one that accurately describes the robotics movement [11]. This work proposes an alternative approach to parameterize the model using the one-step backward (OSBK) data. The resulting deep incremental model serves as a different prediction modality in the MBRL field. The formulated deep incremental model degrades the model learning problem into an easier parametric matrix learning problem, which is especially favourable for high-dimensional robots. We follow the model-accuracy-oriented methods to train the deep incremental model offline.
@@
Model usage and policy learning: When the learned model is ready, the subsequent problem is how to efficiently use the learned model to improve the learning efficiency. The learned model is usually used in a Dyna-style [3, 4] where the learned model is used to generate imagined data for the policy learning process. Note that the imagined data is generated without interacting with the environment. Besides, the learned model is used to compute explicit gradient information [5, 6], which offers inferences for policy learning. Additionally, the works [7, 8] use the learned model to conduct planning for short rollouts. We follow the Dyna-style to get additional training data for the policy learning process.
3 Background Material
@@
This paper considers a Markov decision process (MDP) represented by $\mathcal{M}=\left(\mathcal{S},\mathcal{A},f,r,\gamma\right)$, where $\mathcal{S}\in\mathbb{R}^{n}$, $\mathcal{A}\in\mathbb{R}^{m}$ are the state and action spaces, respectively; $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function; $\gamma\in\left(0,1\right)$ is the discount factor. Denoting $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$ as the robotic state and action. The robotic movement would be described by the transition function $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ that follows
@@
The continuous function $f$ is often unknown. The robot in the state $s_{t}$ interacts with its surrounding environment by applying the action $a_{t}$ according to the policy $\pi(a|s)$, and evolves into the next state $s_{t+1}$ and obtains the reward $r(s,a)$. In the following, we show how to learn a latent-space model $\hat{f}$ in Section 4 and how to use the learned model $\hat{f}$ to improve the learning efficiency in Section 5.
@@ 4 Deep Incremental Model
This section first utilises the OSBK data and the control-theoretical knowledge to transform the nonlinear transfer function (1) into an equivalent parameterized linear form, whose parameters are then learned from input-out data in an offline way
@@ 4.1 Incremental Model Construction
This work introduces a constant matrix $L$ to facilitate the following model parameterization. We first rewrite the robotic evolution model (1) as
$s_{t+1}=La_{t}+h_{t},$ (2)
@@
$s_{t+1}=s_{t}+(L+\frac{h_{t}-h_{t-1}}{\Delta a_{t}})\Delta a_{t}=s_{t}+L_{t}\Delta a_{t},$ (4)
@@
Remark 2.
The observed benefits of delayed data for training is consistent with the results reported in [peng2018sim]. Previous works have empirically found the effectiveness of the delayed data for training robotic policies[peng2018sim], wherein the previous one-step backward action is appended to the original state to construct the augmented state vector for training.
4.2 Model Learning
This subsection uses the approximation ability of the deep neural network (DNN) to learn the parametric matrix $L_{t}$ by minimizing the difference between the true and the predicted next-step values.
This work first represents $L_{t}$ in (4) with a fully connected, multi-layer neural network. Then, the Adam gradient ascent algorithm is adopted to train the DNN on the dataset $\mathcal{D}$ via minimizing the model prediction error :
$L(\theta):=\operatorname*{arg\,min}_{L(\theta)}\sum_{i=1}^{n}\left\|s_{t,i}-(s_{t-1,i}+L(\theta)\Delta a_{t,i})\right\|,$ (5)
Finally, we get the learned incremental evolution model
@@ 4.2 Model Learning
which is the learned representation of (1). The learned deep incremental momodel is in a physics-informed form, wherein the system response within two successive steps is utilized to facilitate learning. This departs from approaches that directly utilizes a DNN to attempt to model the mapping between inputs and outputs. The following section will discuss how to use the learned incremental evolution model (6) to improve the learning efficiency.
@@
We could learn the deep incremental model offline using precollected data, or learn the model along with the policy learning process. For the online model learning scenario, the data is scarce to learn a satisfactory prediction model at the beginning of the model learning process. We mitigate this problem in practice by choosing an initial value for $L(\theta)$ using the prior knowledge of the evolution model (1) if available. The chosen initial value serves as the learning starting point of the $L(\theta)$ learning process. For example, the prior mass matrix of the robot manipulator, although not accurate, is enough to act as the initial value of the $L(\theta)$ learning process.
@@ 5 Policy Learning with Learned Model
This section presents how to use the learned model to improve learning efficiency and how to get the robotic policy.
@@ 5.1 Model Usage
This subsection uses the learned deep incremental model (6) to conduct one-step forward prediction. These predictions and the collected environmental data together serve as sample data for policy evaluation and improvement. We use one-step backward data informed model for one-step forward prediction. This contributes to fully utilizing the robotic physical information to improve learning efficiency.
@@ 5.2 Policy Optimization
This work focuses on the continuous robotic control problem. Therefore, we choose the soft actor-critic (SAC) algorithm to solve our problem. In SAC, the critic agent estimates
$Q^{\pi}(s,a)=E_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right],$ (7)
using the Bellman backup operator, and the actor finds the policy $\pi$ via minimizing the expected KL-divergence
$J_{\pi}(\phi,\mathcal{D})=\mathbb{E}_{s_{t}\sim\mathcal{D}}\left[D_{KL}(\pi||\exp\left\{Q^{\pi}-V^{\pi}\right\})\right].$ (8)
The corresponding pseudo code is provided in Algorithm 1.
@@
Note that our learned deep incremental model (6) is agnostic to the policy learning algorithm. This implies the generality of the MBRL framework in this work. The model learning and usage modules could be used together with different policy optimization algorithms.
Algorithm 1 Deep Incremental Model-Based Reinforcement Learning
@@
3:     Train $f(\theta)$ with $\mathcal{D}_{\text{env}}$ via maximum likelihood
@@
8:              Perform one-step model rollout starting from the sample using policy and add to $\mathcal{D}_{\text{model}}$
@@
6 Online Fine-tuning
The robotic policy trained in simulators might perform undesirably on real robots due to the gap between simulators and hardware. In addition, the policy trained in simulators might not be fully executed, as the trained policy is often directly clipped to satisfy hardware constraints. The above concerns motivate us to fine-tune the incremental policy for enhanced performance online.
Applying the incremental policy $\Delta a_{d}$ (trained in the simulator) to real robots yields the error $e_{k}:=s_{k}-s_{d}$, whose evolution could be described by
$e_{k+1}=e_{k}+L(\theta)\Delta a_{e,k},$ (9)
wherein $\Delta a_{e}$ is the residual incremental policy to be learned online to refine the pretrained $\Delta a_{d}$.
The error evolution model (9) is in a linear form, which is convenient to use model predictive control or linear quadratic regulator to design the residual incremental policy $\Delta a_{e}$ that fine-tunes the pre-trained incremental policy $\Delta a_{d}$ to accommodate the ever-changing environment.
Remark 6.
The deep incremental model learned offline makes the online fine-tuning incremental policy possible. Here the tools from the learning and control fields collaborate to offer a robotic policy with enhanced performance.
@@ 7 Numerical Simulation
The MBPO work inspires this work. This work follows the MBPO work’s way to use the learned model and to optimize a policy. Through comparison with this milestone work, we check whether our modelling technique is more efficient than other modelling techniques. The comparison results presented in Figure 2 show that our approach achieves comparable performance with the MBPO. However, the modelling technique in our approach holds the potential to apply to high-dimensional robots, while the Gaussian process modelling technique used in MBPO is restricted to the low-dimensional domain.
@@ 8 Discussion
This work jointly learns and improves the model and policy from environmental interactions. The control-theoretical knowledge and the OSBK data represent the robotic evolution model as one linear incremental form that contributes to efficient model learning. The learned deep incremental model serves as the prediction model in MBRL and improves the sampling efficiency. The formulated deep incremental model serves as one promising alternative modelling technique in MBRL. The comparative numerical simulations on benchmark continuous robotics control tasks validate the efficiency of our approach.
Limitations The limitation is that how much nonlinearity the deep incremental model could be addressed remains to be clarified. Besides, the deep incremental model is hard to represent discontinuous dynamic systems.
@@ 9 Future Works
The utilized control-theoretical knowledge benefits RL with enhanced sample efficiency by offering an explicit control-oriented deep incremental model, rather than just a black-box input-output mapping. Given stability analysis and safety checks both require one mathematical form of robotic evolutions, the explicit learned model offers us avenues to address safety and stability concerns, which remain to be further explored. This work uses the learned model to generate additional training data. This is one of the ways to utilize the learned model. It is interesting to explore the different roles of the learned model in the learning process. For example, the learned incremental evolution model serves as a basis for planning. Besides, it is also interesting to investigate the performance of utilizing the deep incremental model in an ensemble way.